# Very Basic Set theory

### v7

@@ 5. Axiom ‘pairs’
Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. Otherwise $v$ is not an individual and $w$ and $v$ only have individuals in common.
@@ 6. Axiom ‘union’
The members of the members of a set constitute a set. We postulate ‘union’.
@@
Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
@@
such that $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
Set $z$ is called ‘the transitive closure’ of $s$. $\Box$
The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘set of all natural numbers’.
@@ 7. Natural numbers
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, it is not excluded to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’. Then $\bigcup\omega=\omega$. Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers with first number $\alpha$. Therefore the logic in Very basic set theory is capable to deal with transfinite numbers.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first. There is no ‘empty set’ and regularity is not an axiom but a theorem.
There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$
is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
@@
members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold. $\Box$

### v8

@@ 5. Axiom ‘pairs’
Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. And if $\exists v(v\notin v\wedge v\in w)$ then $\exists v(v\notin v\wedge v\in w)$ such that $v$ and $w$ only have individuals in common.
@@ 6. Axiom ‘union’
The constituents of the members of a set constitute a set. We postulate ‘union’.
@@
Thus there is a rule to construct for every set $z\neq s$ belonging to the sequence a
series of sets $t$, $u$, $v$, , $x$, $y$, $z$ such that each term is a member of the previous
one: $t\in s$, $u\in t$, $v\in u$, , $x\in v$, $y\in x$, $z\in y$ contradictory to regularity if
the sequence is ‘infinite’: the naming of individuals is logically first.
@@
such that $\bigcup w=w$. Then $s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ is a set, and is called ‘the
transitive closure of $s$’. $\Box$
The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof requires no other axioms or no ‘set of all natural numbers’.
@@ 7. Natural numbers
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, it is not prohibited to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’. Then $\bigcup\omega=\omega$. Therefore we can define a number sequence with first number $\omega$ analogous to natural numbers with first number $\alpha$. Then there are the finite ordinals (natural numbers) with first number $\alpha$, followed by ‘transfinite ordinal numbers with first number $\omega$’. And if we define $\omega+\omega=\omega\cdot 2$ then $\bigcup(\omega\cdot 2)=\omega\cdot 2$ thus the next sequence of transfinite ordinals are the ordinals with first number $\omega\cdot 2$, and so on, that is
$\alpha$, $\alpha+1$, $\alpha+2$, …, $\omega$, $\omega+1$, $\omega+2$, …, $\omega\cdot 2$, $\omega\cdot 2+1$, $\omega\cdot 2+2$, …, $\omega\cdot 3$, ….
Thus the logic in Very basic set theory is capable to deal with transfinite numbers.
@@ Conclusion
Very basic set theory is a set theory with the philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals. The atomic facts are the membership relations and the elementary propositions are the individuals. The formulas of set theory preserve definiteness. The naming of individuals is logically first and the logic is capable to deal with transfinite numbers.
We list some differences with ‘mainstream set theory’.
  • There is no ‘empty set’ (Kanamori [2003] §1).
  • The singleton of an individual is that individual itself (Kanamori [2003] §2).
  • Regularity is not an axiom but a theorem.
  • The proof that ‘every set has a transitive closure’ requires no other axioms or no ‘set of all natural numbers’ (Mendelson [2015] pp.286-287).
@@
    members’ of the powerset ${\mathcal{P}}s$.
    However, there is no ‘empty set’ therefore ${\mathcal{P}}\{s\}=\{\{s\}\}$.
    Thus Cantor’s Theorem does not hold.
  • Any pair of individuals is the first number of a sequence of natural numbers. Then transfinite ordinals are postulated. If e.g. $\bigcup\omega=\omega$ then $\omega$ is the first number of a sequence of transfinite ordinals. Therefore $\omega$ is not a ‘limit ordinal’, $\omega$ is the first number of a sequence satifying the Peano Axioms for natural numbers.